ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 76 Dziri, Raja, Faculté des Sciences de Tuni, Tunisi Dynamical shape optimaization in Navier-Stokes equations. We study a dynamical shape optimisation problem in a non-cylindrical Navier-Stokes flow. The purpose is to determine the optimal tube descirbed by the flow during a fiven finite time-interval in order to bring it to rest. It consists in searching the non-autonomous vector field that builts such tube. The particularity of that problem is the necessarity to introduce two adjoint states to be able to derive the optimal necessary condition. This work is jointly done with Prof. Jean Paul Zolesio, France.